OpenStudy (anonymous):

What is the product of the roots of the equation x^2-3x+7=0?

OpenStudy (anonymous):

the old good equation says that x1.x2=c/a so it is 7/1=7

OpenStudy (anonymous):

If you have: \[x^2+bx+c=0\] it is just c.

OpenStudy (anonymous):

yeah! that's right: c/1=c